A project has an initial cost of $38,550, expected net cash inflows of $10,000 per year for 8 years, and a cost of capital of 10%. What is the project's MIRR? (Hint: Begin by constructing a time line.) Do not round intermediate calculations. Round your answer to two decimal places.

*Please show how to calculate MIRR with BA II Plus*
Answer #1

Initial Cost = $ 38500, Expected Net Cash Inflows = $ 10000, Tenure = 8 years and Interest Rate = Cost of Capital = 10 %

Solving the problem using a BA II Plus:

- Input: PMT = $ 10000, N = 8 years , I/Y = 10 %, PV = 0

- CMPT -> FV

- FV = $ 114358.88

- PV = Initial Cost = $ 38500

- MIRR = [FV/PV]^(1/N) - 1 = [114358.88 / 38500]^(1/8) - 1 = 0.14578 or 14.578 % ~ 14.58 %